### v3

# Detect Everything with Few Examples

Abstract
Few-shot object detection aims at detecting novel categories given a few example images. Recent methods focus on finetuning strategies, with complicated procedures that prohibit a wider application. In this paper, we introduce DE-ViT , a few-shot object detector without the need for finetuning. DE-ViT ’s novel architecture is based on a new region-propagation mechanism for localization. The propagated region masks are transformed into bounding boxes through a learnable spatial integral layer. Instead of training prototype classifiers, we propose to use prototypes to project ViT features into a subspace that is robust to overfitting on base classes. We evaluate DE-ViT on few-shot, and one-shot object detection benchmarks with Pascal VOC, COCO, and LVIS. DE-ViT establishes new state-of-the-art results on all benchmarks. Notably, for COCO, DE-ViT surpasses the few-shot SoTA by 15 mAP on 10-shot and 7.2 mAP on 30-shot and one-shot SoTA by 2.8 AP50. For LVIS, DE-ViT outperforms few-shot SoTA by 20 box APr.
Preprint.
1 Introduction
Object recognition and localization are two of the core tasks in computer vision. Few-shot object detection provides a promising paradigm for generic object detectors by representing novel categories with a set of support images [2]. However, despite the promising practicality in principle, most recent few-shot methods rely on finetuning to adapt to novel classes  [28], which is often a complicated and tedious procedure that limits the practical use of these methods and that results in a large accuracy gap between base and novel classes [55].
Motivated by the recent progress in unsupervised object discovery with pretrained vision transformers (ViTs) [36, 43], strong pretrained ViTs can also be used to overcome the limitations of finetuning. However, despite their rich semantical representations, pretrained ViT features lack the coordinates information that is required to perform a bounding box regression. As shown in Sec. 4.2, naively applying a conventional regression on ViT features yields poor localization results, while unfreezing the ViT backbone leads to an accuracy collapse on novel classes, by completely overfitting the base classes.
To address these challenges, we propose a novel region-propagation-based localization architecture. In the proposed architecture, object proposals are expanded by a fixed ratio. Objects are localized by performing a mask prediction within the expanded proposals instead of a bounding box regression. To accurately derive bounding boxes from the propagated regions, we propose the spatial integral layer, a learnable mask-to-box transformation. To further narrow the accuracy gap between base and novel classes, we propose to construct prototypes not as classifier weights, as shown in Fig. 2(a), but to project ViT features into a subspace that is used as network inputs. Our empirical studies demonstrate that the projected features are more robust to overfitting on base classes.
Figure 1: Demonstration of the proposed method DE-ViT on YCB objects [4]. Note that the model is trained on only the base categories of LVIS. A few example images of YCB objects are provided as novel categories during inference only.
With these proposed techniques, we introduce DE-ViT , a few-shot object detector that uses example images to detect novel objects without the need for any finetuning. An example of the results returned by DE-ViT is shown in Fig. 1.
We evaluate DE-ViT on few-shot, and one-shot object detection benchmarks with Pascal VOC [9], COCO [31] and LVIS [15] datasets. Our method establishes new state-of-the-art (SoTA) results on all benchmarks. For COCO, DE-ViT surpasses the SoTA LVC [26] by 15 mAP on 10-shot and by 7.2 mAP on 30-shot, and it also surpasses the one-shot SoTA BHRL [52] by 2.8 AP50. For Pascal VOC, DE-ViT surpasses the SoTA NIFF [14] by 2.0 nAP50. For LVIS, which has been regarded as a highly challenging dataset for few-shot object detection [46], DE-ViT outperforms the SoTA DiGeo [34] by 20 box APr. Notably, our method achieves a faster inference time while having a better accuracy.
Our contributions are summarized as follows. (1) We introduce DE-ViT , a fast few-shot detector that detects novel objects without any finetuning. (2) We propose a novel region-propagation-based architecture for object detection, a spatial integral layer that learns to transform masks into boxes, and a new feature projection that significantly reduces overfitting on base classes. (3) We demonstrate that DE-ViT establishes state-of-the-art performance over few-shot and one-shot benchmarks on Pascal VOC, COCO, and LVIS.
2 Related work
Figure 2: Existing meta-learning-based FSOD methods can be divided into two categories (a and b). Methods in the first category (a) build prototypes from novel class examples and use these prototypes as the classifier weights of a detection network. Despite its simplicity, this strategy exhibits inferior accuracy [28]. Methods in the second category (b) learn to match the proposal regions in the query image and novel examples through a matching network. This strategy is computationally heavy due to dense feature interactions across multiple images and usually requires finetuning to increase accuracy in novel classes [18, 19]. In contrast, our method (c) applies a dot-product with the prototypes to project ViT features into a subspace that is robust to overfitting on base classes, and then applies a region propagation network to refine the localization and derive the class score. Our method does not employ any finetuning for novel classes.
Few-shot Object Detection (FSOD) aims at detecting objects of novel classes by utilizing a few support images from novel classes as training samples [28]. Existing approaches can be broadly classified into finetuning-based [46, 11, 45, 50, 14] and meta-learning-based strategies [51, 25, 10]. Finetuning-based methods, despite their prevalence, suffer from a large accuracy gap between the base and novel classes, as well as practical limitations due to redundant multi-stage procedures [55]. Meta-learning-based methods avoid finetuning through online adaptation. Early works on meta-learning FSOD (Fig. 2.a) construct prototypes from novel class examples and use the prototypes as input to a network that classifies query images [51, 25]. Recent methods (Fig. 2.b) design interaction mechanisms of dense spatial features between query and support images [18, 19, 17]. In contrast, our method (Fig. 2.c) computes dot-products with the prototypes to project their features into a subspace, instead of using a prototype classifier, and applies a region-propagation network instead of a dense feature matching network. One-shot Object Detection (OSOD) is an extreme case of FSOD with only one example per novel class, it reduces the setting to single-class detection [52].
Vision Transformer (ViT) is a recent network architecture that demonstrates stronger performance and more interpretable features than traditional convolutional architectures. There is a growing trend of applying self-supervised ViTs, such as DINO [5, 39], to unsupervised object discovery [54, 59, 41, 43]. TokenCut [47] applies graph-cut over DINO features to separate the primary foreground object. DeepSpectral [36] predicts semantic segmentation masks without supervision through spectral clustering over DINO features. OW-DETR [16] uses DINO ViT to discover unknown objects in an open-world setting. In this work, we use pretrained ViTs for solving the few-shot object detection problem.
3 Method
3.1 Problem Formulation
We use $\mathcal{C}$ to denote the set of classes. In few-shot object detection (FSOD), $\mathcal{C}$ is composed of a set of base classes, denoted by $\mathcal{C}_{base}$, and a set of novel classes, denoted by $\mathcal{C}_{novel}$. Thus, $\mathcal{C}=\mathcal{C}_{base}\cup\mathcal{C}_{novel}$ and $\mathcal{C}_{base}\cap\mathcal{C}_{novel}=\varnothing$. During training, a large number of examples are provided for the base classes. During testing, only $k$ labeled samples are provided for each novel class. The samples for novel classes are referred to as support images. The goal of FSOD is to leverage the training data of $\mathcal{C}_{base}$ to learn a detector that can detect objects of $\mathcal{C}_{novel}$ given the $k$-shot support images.
3.2 Region Propagation Network
Figure 3: Overall framework of our proposed method. Given a proposal box of a query image, we extract the initial mask region (shown here in yellow) and expand the proposal by a constant ratio. Next, the region within the expanded proposal is gradually propagated and refined to fit the object area through a sequence of propagation layers. Each propagation layer accepts previous regions and features as input while returning updated features and region as outputs. The final predicted region is transformed to bounding box coordinates through a learnable spatial integral layer, as detailed in Sec. 3.3. The predicted region also serves as spatial attention and is used for a weighted averaging of the features along height and width. The averaged features are then mapped to class scores. The projected features are the dot products between the ViT features of a given query image and class prototypes, as detailed in Sec. 3.4.
Despite having rich semantics information, pretrained ViT features lack the coordinates information required for bounding box regression. As shown in Sec. 4.2, naively applying a conventional regression on ViT features yields poor localization results. A natural solution is to learn this localization capability by finetuning the ViT backbone during the training of the detector with the base classes. However, we observed that finetuning results in completely overfitting the base classes and in an accuracy collapse on novel classes. This was observed when integrating DINOv2 ViT into the framework of Meta RCNN [51], a standard prototype-based FSOD, as shown in Sec. 4.2. This suggests that harnessing the generalization power of strong ViT backbones for FSOD is non-trivial. The question here is how to produce accurate localization with pretrained ViT features.
Given an object proposal, we use a region propagation network that gradually propagates the proposal region to accurately cover and fit the object by refining an object mask. Unlike bounding box regression, mask prediction localizes objects without coordinate outputs. The propagated region is then transformed into a bounding box through a learnable spatial integral layer. We use an off-the-shelf region proposal network (RPN) to generate the initial region proposals, as class-agnostic proposals are shown to generalize well to novel classes [13]. Each proposal is expanded by a fixed ratio in order to delimit the propagation boundaries. The overall framework is shown in Fig. 3.
Figure 4: Visualizations of region propagation on classes “horse” and “birds” (first and second row). The proposal, expanded proposals, and final predicted boxes are colored in transparent yellow, yellow, and green, respectively. The three propagated regions (from left to right) are sampled from the output of the first three PL blocks in ascending order.
Propagation Layer. We propose the region Propagation Layer (PL), a new type of network module designed for object detection. PL serves as the central building block of our method. An example is shown in Fig. 4. The $t$-th PL block takes all previous regions $r_{0:t-1}\in\mathbb{R}^{t\times K\times K}$ and the previous PL block features $h_{t-1}\in\mathbb{R}^{S\times K\times K}$ as input, where $t$ denotes the number of PL blocks, $S$ denotes the number of feature channels, and $K$ denotes the feature spatial size. The $t$-th PL block outputs the updated region $r_{t}\in\mathbb{R}^{1\times K\times K}$, features $h_{t}\in\mathbb{R}^{S\times K\times K}$, bounding box $b_{t}\in\mathbb{R}^{4}$, and class score $c_{t}\in\mathbb{R}$. Each PL block works as a small detection network and can be stacked to improve accuracy. The update rule is explained in Eq. 1 and illustrated in the third part of Fig. 3.
$h_{t}=f_{\text{update},t}(\operatorname{concat}(r_{0:t-1},h_{t-1});\theta),\,\,h_{t,region}=f_{\text{region},t}(h_{t};\theta)$ (1)
$r_{t}=\sigma(h_{t,region}),\,\,b_{t}=f_{\text{integral},t}(h_{t,region};\theta)$
$c_{t}=f_{\text{class},t}(\text{\scriptsize WeightedAvgPool}(h_{t},r_{t});\theta)$
In Eq. 1, $f_{\text{update},t}$ denotes the conv-bn-relu block with $S$ output channels that updates the hidden features, and concat denotes channel-wise concatenation. $f_{\text{region},t}$ denotes the conv block with single channel output that predicts the output region logits $h_{t,region}\in\mathbb{R}^{1\times K\times K}$. $\sigma$ denotes the sigmoid function. $f_{\text{integral},t}$ denotes the spatial integral layer detailed in Sec. 3.3. $h_{t}$ is aggregated over spatial dimensions to $\text{\footnotesize WeightedAvgPool}(h_{t},r_{t})\in\mathbb{R}^{S\times 1\times 1}$ with weights $r_{t}$. $f_{\text{class},t}$ denotes the linear block that maps $\text{\footnotesize WeightedAvgPool}(h_{t},r_{t})$ to class scores. $\theta$ denotes network parameters. During training, we use focal loss and L1 regression loss for the output class score $c_{t}$ and bounding box $b_{t}$. For the output region $r_{t}$, we apply BCE loss and Dice loss [44]. The region labels during training are generated by the ground-truth object region within the expanded proposals.
3.3 Learnable Spatial Integral
Converting masks to bounding boxes is a widely-used transformation in instance segmentation networks [40] as a post-processing step. The standard solution is to find the top-left and bottom-right foreground pixels and use their positions as the bounding box coordinates [35]. However, this approach has major limitations. Firstly, this mask-to-box conversion assumes the availability of ground-truth instance masks, which are much more expensive to obtain than bounding boxes [31]. Moreover, this approach is non-differentiable and is also prone to outliers. Therefore, the question is how to accurately derive bounding boxes from the region-based localization results using a learnable and differentiable function.
Let $b^{\text{out}}=(c_{w}^{\text{out}},c_{h}^{\text{out}},w^{\text{out}},h^{\text{out}})$ denote the output bounding box, where $c_{w}^{\text{out}}\in[0,W],w^{\text{out}}\in[0,W]$ and $c_{h}^{\text{out}}\in[0,H],h^{\text{out}}\in[0,H]$. Instead of predicting $b^{\text{out}}$ directly, we propose to first predict a relative bounding box $b^{\text{rel}}=(c_{w}^{\text{rel}},c_{h}^{\text{rel}},w^{\text{rel}},h^{\text{rel}})\in[0,1]^{4}$, that can be transformed to $b^{\text{out}}$ according to Eq. 2,
$\begin{split}&(w^{\text{out}},h^{\text{out}})=(w^{\text{exp}}\,w^{\text{rel}},h^{\text{exp}}\,h^{\text{rel}}),\\ &(c^{\text{out}}_{w},c^{\text{out}}_{h})=(c_{w}^{\text{exp}}-0.5w^{\text{exp}},c_{h}^{\text{exp}}-0.5h^{\text{exp}})+(c^{\text{rel}}_{w}\,w^{\text{exp}},c^{\text{rel}}_{h}\,h^{\text{exp}}),\end{split}$ (2)
where $b^{\text{exp}}=(c_{w}^{\text{exp}},c_{h}^{\text{exp}},w^{\text{exp}},h^{\text{exp}})$ denotes the expanded proposal. Thus, $b^{\text{rel}}$ is a normalized bounding box relative to $b^{\text{exp}}$. Let $h_{region}\in\mathbb{R}^{K\times K}$ denote the output region logits, where we skip the notation of $t$-th block and the channel of $1$ for simplicity. Our spatial integral layer $f_{\text{integral}}$ estimates $b^{\text{rel}}$ with Eq. 3 and 4. An illustrative example is given in Fig. 5.
[Uncaptioned image] Figure 5: The max row sum and column sum are used as the box width and height of the triangular object.
[…]
$w^{\text{rel}}=\sum_{i=1}^{K}\sum_{j=1}^{K}\frac{\sigma(h_{region})_{(i)j}}{K}\mathbf{\theta^{w}}_{i}$ (4)
$h^{\text{rel}}=\sum_{j=1}^{K}\sum_{i=1}^{K}\frac{\sigma(h_{region})_{i(j)}}{K}\mathbf{\theta^{h}}_{j}$
To motivate Eq. 3 and 4, consider the toy example of converting a binary triangle mask to a bounding box in Fig. 5. A reasonable approach is to compute the mask center as the bounding box center and use the max row sum and column sum as width and height. Inspired by this insight, we compute the expected position of the spatial distribution $\text{softmax}(h_{region})$ as the center of $b^{\text{rel}}$ in Eq. 3. We compute the row and column sums of the output region as $\sum_{j=1}^{K}\sigma(h_{region})_{ij}$ and $\sum_{i=1}^{K}\sigma(h_{region})_{ij}$ in Eq. 4. Instead of picking the maximum, we apply a soft aggregation to average all row or column sums in terms of magnitude rank. The aggregation is done by sorting the row and column sums, and then computing the weighted average. This explains the use of order statistics notation $(i)$ and $(j)$. Parameters $\mathbf{\theta^{h}}\in\mathbb{R}^{K},\mathbf{\theta^{w}}\in\mathbb{R}^{K}$ are learnable aggregation weights.
3.4 Feature Subspace Projection
A major challenge of FSOD is to generalize to novel classes that are unseen during training. However, despite numerous attempts to solving this problem, by using margin-based regularization for example [34], there persists a considerable accuracy gap between base and novel classes [2]. This disparity indicates that a network trained with base classes would inevitably overfit on patterns that are only present among the base classes. A classic technique to reduce overfitting consists in representing data in a low-rank subspace [42]. We explore in this work the construction of a subspace of pre-trained ViT features that reduces the accuracy gap between base and novel classes.
Prototypes are class representatives built from support images. Given the support images of each class, we compute the ViT features, crop the features with object bounding boxes and use the average feature as the class prototype [51]. Let $p_{\mathcal{C}}\in\mathbb{R}^{|\mathcal{C}|\times D}$ denote the prototypes of classes from set $\mathcal{C}$, where $D$ denotes the channel dimension. Let $h_{vit}\in\mathbb{R}^{D\times H\times W}$ denote the ViT features of the query image. We assume that both prototypes and features are normalized to unit length at the channel dimension. Then $p_{\mathcal{C}}\cdot h_{vit}\in\mathbb{R}^{|\mathcal{C}|\times H\times W}$ can be interpreted as a subspace projection with $p_{\mathcal{C}}$ being the basis. However, this subspace construction has two limitations. Firstly, only using prototypes of classes set $\mathcal{C}$ can be too limited to sufficiently capture the feature information. Secondly, a permutation of $\mathcal{C}$ creates a different but equivalent subspace, yet designing permutation-invariant networks is a highly challenging problem [22]. For the first limitation, we introduce a set $\mathcal{B}$ of background classes, $\mathcal{B}\cap\mathcal{C}=\varnothing$, with $p_{\mathcal{B}}\in\mathbb{R}^{|\mathcal{B}|\times D}$ being the prototypes of $\mathcal{B}$, to preserve more information from $h_{vit}$. For the second limitation, we propose to build a separate subspace for each class $c\in\mathcal{C}$, and reorder other classes $\mathcal{C}\setminus c$ to resolve permutation ambiguity. The feature subspace projection is explained in Eq. 5 and illustrated in Fig. 3.
$h_{subspace,c}=\operatorname{concat}(p_{c}\cdot h_{vit},\operatorname{channel-reorder}(p_{\mathcal{C}\setminus c}\cdot h_{vit}),p_{\mathcal{B}}\cdot h_{vit})$ (5)
In Eq. 5, $h_{subspace,c}\in\mathbb{R}^{S\times H\times W}$ denotes the subspace feature for class $c$, and function $\operatorname{channel-reorder}$ sorts channels by magnitude and then interpolates $|\mathcal{C}|-1$ channels to a pre-defined number of channels. In practice, we use example images of non-object stuff classes, e.g., sky, road, floor, to construct $\mathcal{B}$. As shown in Sec. 4.2, feature subspace projection significantly reduces the accuracy gap between base and novel classes. On the other hand, creating a subspace for each class $c\in\mathcal{C}$ introduces costly per-class inference. However, the per-class inference cost can be reduced by finding the top $T$ most likely classes with a lightweight prototype classifier [51] and only performing inference for these $T$ classes. As shown in Sec. 4.2, our method achieves a faster inference speed and surpasses SoTA when $T=3$.
Table 1: Results on COCO 2014 few-shot benchmark. Our method outperforms existing work in detecting novel classes by a significant margin. Results surpassing the SoTA are indicated in bold.
[…]
FSRW [25] - 5.6 12.3 4.6 - 9.1 19 7.6
Meta R-CNN [51] 5.2 6.1 19.1 6.6 7.1 9.9 25.3 10.8
TFA [46] 33.9 10 19.2 9.2 34.5 13.5 24.9 13.2
Multi-Relation Det [10] - 16.6 31.3 16.1 - - - -
FSCE [45] - 11.9 - 10.5 - 16.4 - 16.2
Retentive RCNN [11] 39.2 10.5 19.5 9.3 39.3 13.8 22.9 13.8
HeteroGraph [17] - 11.6 23.9 9.8 - 16.5 31.9 15.5
FsDetView [50] 6.4 7.6 - - 9.3 12 - -
Meta Faster RCNN [18] - 12.7 25.7 10.8 - 16.6 31.8 15.8
LVC [26] 28.7 19 34.1 19 34.8 26.8 45.8 27.5
CrossTransformer [19] - 17.1 30.2 17 - 21.4 35.5 22.1
NIFF [14] 39 18.8 - - 39 20.9 - -
DiGeo [34] 39.2 10.3 18.7 9.9 39.4 14.2 26.2 14.8
DE-ViT (Ours) ViT-S/14 24 27.1 43.1 28.4 24.2 26.9 43.1 28.5
ViT-B/14 28.3 33.2 51.4 35.5 28.5 33.4 51.4 35.7
ViT-L/14 29.4 34.0 52.9 37.0 29.5 34.0 53.0 37.2
4 Experiments
We comprehensively evaluate our method on few-shot and one-shot benchmarks. Furthermore, we compare the efficiency of our method against SoTA solutions, study few-shot performance for different numbers of shots, compare it to language-based detectors, and provide qualitative results. We conduct ablations to study the contributions of the proposed components to the performance of our method. Our source code and the pretrained models will be publicly released upon acceptance.
Evaluation Metrics and Datasets. Few-shot and one-shot evaluations split classes into base and novel classes. Base classes are seen during training and novel classes are unseen. The performance on novel classes is more important. For COCO and Pascal VOC, nAP, nAP50, and nAP75 represent mAP, AP50, and AP75 in novel classes. bAP and bAP50 represent mAP and AP50 in base classes. One-shot evaluation conventionally divides 80 classes of COCO into four even partitions, and alternatively uses three as base classes and one partition as novel classes [37]. There are 4 base/novel splits in total, named Split-1/2/3/4. For LVIS, APr, APc, and APf represent AP on rare, common, and frequent categories. Rare categories are used as novel classes. Metrics on LVIS, such as box APr and mask APr, are computed on bounding boxes and on instance segmentation masks separately. We evaluate our method on Pascal VOC [9], COCO [31], and LVIS-v1 [15]. We follow the conventional base/novel classes split and use the same support images of novel classes with existing work [46, 52].
Table 2: nAP50 results on Pascal VOC few-shot benchmark. Results surpassing the SoTA are indicated in bold. (*) denotes that implementation is not publicly available.
Method Novel Split 1 Novel Split 2 Novel Split 3 Avg
1 2 3 5 10 1 2 3 5 10 1 2 3 5 10
TFA [46] 39.8 36.1 44.7 55.7 56.0 23.5 26.9 34.1 35.1 39.1 30.8 34.8 42.8 49.5 49.8 39.9
FsDetView 25.4 20.4 37.4 36.1 42.3 22.9 21.7 22.6 25.6 29.2 32.4 19 29.8 33.2 39.8 29.2
Multi-Relation Det [10] 37.8 43.6 51.6 56.5 58.6 22.5 30.6 40.7 43.1 47.6 31 37.9 43.7 51.3 49.8 43.1
Retentive RCNN [11] 42.4 45.8 45.9 53.7 56.1 21.7 27.8 35.2 37.0 40.3 30.2 37.6 43 49.7 50.1 41.1
Meta Faster R-CNN [18] 43.0 54.5 60.6 66.1 65.4 27.7 35.5 46.1 47.8 51.4 40.6 46.4 53.4 59.9 58.6 50.5
LVC [26] 54.5 53.2 58.8 63.2 65.7 32.8 29.2 50.7 49.8 50.6 48.4 52.7 55 59.6 59.6 52.3
CrossTransformer [19] 49.9 57.1 57.9 63.2 67.1 27.6 34.5 43.7 49.2 51.2 39.5 54.7 52.3 57 58.7 50.9
HeteroGraph [17] 42.4 51.9 55.7 62.6 63.4 25.9 37.8 46.6 48.9 51.1 35.2 42.9 47.8 54.8 53.5 48.0
DiGeo [34] 37.9 39.4 48.5 58.6 61.5 26.6 28.9 41.9 42.1 49.1 30.4 40.1 46.9 52.7 54.7 44.0
NIFF [14] (*) 62.8 67.2 68.0 70.3 68.8 38.4 42.9 54.0 56.4 54 56.4 62.1 61.2 64.1 63.9 59.4
DE-ViT (Ours) ViT-S/14 47.5 64.5 57.0 68.5 67.3 43.1 34.1 49.7 56.7 60.8 52.5 62.1 60.7 61.4 64.5 56.7
ViT-B/14 56.9 61.8 68.0 73.9 72.8 45.3 47.3 58.2 59.8 60.6 58.6 62.3 62.7 64.6 67.8 61.4
ViT-L/14 55.4 56.1 68.1 70.9 71.9 43.0 39.3 58.1 61.6 63.1 58.2 64 61.3 64.2 67.3 60.2
Figure 6: Qualitative visualization of our method DE-ViT on COCO (a) and LVIS (b), and comparison to Meta-Faster RCNN [18]. Boxes of base and novel classes are colored in green and yellow, respectively.
Implementation Details. We adopt a standard two-stage object detection framework, similarly to Mask R-CNN [20]. We use the same off-the-shelf RPN with existing work [56] to generate object proposals. The RPN is trained separately for each dataset using only base classes. We use DINOv2 [39] ViT as the backbone unless specified, and report results in ViT-S/B/L (small, base, large) model sizes. The ViT backbones are kept frozen during detector training. We adopt the prototype extraction procedure of Meta RCNN [51] using ViT features. Prototypes of background classes set $\mathcal{B}$ are extracted from images of stuff (non-object) classes, e.g., sky, road, from COCOStuff [3] unless specified. Similar to [51], the top $T$ most likely classes for each proposal are determined by the distances between prototypes and the average proposal feature. We set $T$ to 10 unless specified, where $T$ is the number of classes to create subspace features and perform inference as explained in Sec. 3.4. We apply 3 PL blocks for experiments on Pascal VOC and COCO, and 5 PL blocks for those on LVIS.
4.1 Main Results
Tab. 3.4 shows our results on the few-shot COCO benchmark. Our method DE-ViT outperforms the previous SoTA LVC [26] by a significant margin (+15 nAP on 10-shot, +7.2 nAP on 30-shot). It is worth noting that LVC requires over ten stages for self-training and pseudo-labeling procedures on novel classes [1], while our method DE-ViT is trained once on the base classes and used directly on the novel classes. A pretrained model for LVC has never been released. We plot the nAP50 of our method and the SoTAs with different numbers of shots in Fig. LABEL:fig:shots.
Tab. 2 shows our results on the few-shot Pascal VOC benchmark. Our method DE-ViT outperforms the previous SoTA NIFF [14] by +2.0 on averaged nAP50. It is worth noting that the implementation of NIFF has not been released.
LVIS has been regarded as a highly challenging benchmark for FSOD [46] with 337 novel classes and only DiGeo [34] reports few-shot results on LVIS v1. Tab. LABEL:tab:lvis-fewshot shows that our method outperforms DiGeo in all metrics and a significant boost in the accuracy of detecting novel objects (+20 box APr).
Tab. 4 shows our results on the one-shot COCO benchmark. DE-ViT surpasses the previous SoTA BHRL by 6 bAP50 and 2.8 nAP50. One-shot object detection task follows a single-class detection setting. Therefore, we adapt our method DE-ViT by detecting each class separately during evaluation. OWL-ViT [38] also reports one-shot results on COCO. However, OWL-ViT’s results are obtained with an ensemble of language-based detection and one-shot pipelines without providing an implementation or isolated measurements. Therefore, we exclude OWL-ViT from the one-shot comparison.
Table 4: Results on COCO 2017 one-shot benchmark. DE-ViT outperforms existing work and is not limited to single class detection and single support image as other one-shot methods.
[…]
Split-1 Split-2 Split-3 Split-4 Avg Split-1 Split-2 Split-3 Split-4 Avg
SiamMask [37] 38.9 37.1 37.8 36.6 37.6 15.3 17.6 17.4 17 16.8
CoAE [23] 42.2 40.2 39.9 41.3 40.9 23.4 23.6 20.5 20.4 22
AIT [6] 50.1 47.2 45.8 46.9 47.5 26 26.4 22.3 22.6 24.3
SaFT [55] 49.2 47.2 47.9 49 48.3 27.8 27.6 21 23 24.9
BHRL [52] 56 52.1 52.6 53.4 53.6 26.1 29 22.7 24.5 25.6
DE-ViT (Ours, ViT-L/14) 59.4 57.0 61.3 60.7 59.6 27.4 33.2 27.1 26.1 28.4
4.2 Analysis and Ablation Studies
Efficiency. We compare the inference time of DE-ViT with different values of $T$ against recent few-shot works in Tab. 5 in COCO 10-shots setting, where Swin denotes Swin Transformer [32], RN101 denotes ResNet101 [21], MFRCNN denotes Meta Faster RCNN [18], and CrossT denotes Cross Transformer [19]. $T$ is the number of classes to create subspace features and perform inference. Tab. 5 shows that our method DE-ViT can achieve the smallest inference time while having better accuracy. The inference times of all the compared methods are measured on the same machine.
Table 5: Ablation study on different values of $T$ and inference time comparison with existing methods. N/A: The pretrained model is not publicly available for evaluation.
Method Backbone $T$ nAP50 Secs/Img
DE-ViT (Ours) ViT-L/14 1 49.6 0.22
3 52.5 0.33
10 52.9 0.83
LVC [26] Swin-S - 34.1 N/A
MFRCNN [18] RN101 - 25.7 0.61
CrossT [19] Custom - 30.2 3
Qualitative results. Fig. 4 shows examples of region propagations on classes “horse” and “bird”. The regions gradually move from the initial proposal region toward the object regions and are transformed into predicted boxes with our spatial integral layer. Fig. 6 shows a qualitative demonstration of our method on COCO and LVIS, and a comparison to Meta Faster RCNN [18]. Note that we chose Meta Faster RCNN over other existing works because of the availability of its pretrained models. DE-ViT detects more novel objects while having fewer false positives. We provide more visualizations in the supplementary material.
Table 6: Comparison to language-based detectors on LVIS and COCO 2017. $\dagger$: use customized pretrained backbone instead of public ones. “-”: result that was not reported.
Method Backbone Use Extra LVIS COCO
Training Set mask APr box APr nAP50
ViLD [13] EffNet-B7 26.3 27 27.6
RegionCLIP [56] RN50x4 - 22 39.3
OV-DETR [53] ViT-B/32 - 17.4 29.4
Detic [58] RN50 17.8 - 27.8
OWL-ViT [38] ViT-L/14$\dagger$ - 25.6 -
OWL-ViT [38] ViT-L/14$\dagger$ - 31.2 -
CORA+ [49] RN50x4 - 28.1 43.1
Ro-ViT [27] ViT-L/14$\dagger$ 31.4 - 33
Co-Det [33] Swin-B 29.4 - 30.6
F-VLM [27] RN50x64 32.8 - 28.0
DE-ViT (Ours, 30-shots) ViT-S/14 24.2 23.4 39.5
ViT-B/14 28.5 26.8 45.4
ViT-L/14 34.3 33.6 50
Comparison to language-based detectors. Tab. 6 compares our method DE-ViT against language-based detectors on the COCO and LVIS-v1 datasets. Language-based detection, also known as open-vocabulary detection [48], does not assume access to support images on novel classes but requires the knowledge of the novel class names. The class names are then used to discover images of novel classes from a large collection of image-text pairs. Therefore, language-based detectors have the following limitations. First, many objects lack clear names [30], e.g., objects that are specific to certain contexts [24]. Second, the association between visual concepts and language is evolving and not static [8]. In contrast, few-shot object detection (FSOD) does not assume any knowledge of class names, and novel classes are described with support images only. FSOD’s setting is therefore more general and arguably more challenging, and aims to emulate humans’ capability to recognize objects by their consistent appearance.
In COCO, our method DE-ViT outperforms the previous SoTA CORA+ by 6.9 AP50. Our method only trains on COCO while CORA+ uses ImageNet-21K [29] and COCO Captions [7] as additional training data. In LVIS, DE-ViT outperforms the previous SoTA on mask APr (+1.5 over F-VLM) and box APr (+2.4 over OWL-ViT). Note that we report the Co-Det [33] result with Swin-B backbone instead of EVA02-L backbone [12] because the latter includes the base and novel classes of COCO in its training set. Moreover, we observe a high variance in the performance of language-based detectors. F-VLM achieves a mask APr of 32.8 on LVIS but only has 28 nAP50 on COCO. While CORA+ has 43.1 nAP50 on COCO but only a box APr of 28.1 on LVIS. On the contrary, our method DE-ViT outperforms existing solutions on both LVIS and COCO.
Table 7: Ablation studies on the feature subspace projection.
Table 9: Ablation studies on the region propagation-based localization.
o c|ccc|ccccRaw Feature Input Feature Subspace Projection Novel Base
Prototype Classifier       $c$          $\mathcal{B}$    $\mathcal{C}\setminus c$ nAP50 nAP75 bAP50 bAP75
4.5 2.2 48.9 22.5
      ✓ 26.2 9.7 29.3 12
      ✓          ✓ 38.4 23 43.4 26.8
      ✓          ✓    ✓ 39.5 24.1 42.3 25.9
Conventional Expanded Region Propagation Novel Base
Regression Net Proposal Network nAP50 nAP75 bAP50 bAP75
37.7 14.6 46.5 23.9
35.6 12.5 41.3 19.8
39.5 24.1 42.3 25.9
Feature subspace projection. We examine the component effects for feature subspace projection in Tab. 4.2. The first column represents the integration of raw features of DINOv2 ViT into Meta RCNN [51], a standard prototype-based few-shot detector. Tab. 4.2 shows that the model has an accuracy collapse on novel classes with raw ViT feature inputs by completely overfitting base classes. The $c$, $\mathcal{B}$, and $\mathcal{C}\setminus c$ represents different groups of prototypes used in feature projection, as in Eq. 5. By projecting features to $p_{c}$, general detection ability emerges (nAP50: 4.5 $\rightarrow$ 26.2), then significantly improves after adding background classes prototypes $p_{\mathcal{B}}$ (nAP50: 26.2 $\rightarrow$ 38.4), and further improves with $p_{\mathcal{C}\setminus c}$ (nAP50: 38.4 $\rightarrow$ 39.5). Results in Tab. 4.2 and 4.2 are obtained on COCO 2017 with ViT-S/14.
Region propagation network. We study the impacts of region-propagation-based localization in Tab. 4.2. The conventional regression network exhibits poor localization accuracy on novel classes (nAP75: 14.6), and simply expanding proposals even lowers the performance (nAP75: 14.6 $\rightarrow$ 12.5). When using our proposed region-propagation network with the learnable spatial integral layer, the localization accuracy has a significant boost (nAP75: 12.5 $\rightarrow$ 24.1). Note that the results in Tab. 4.2 are obtained by only changing the localization architectures while keeping the feature subspace projection.
Number of propagation layers. We study the impacts of the number of PL blocks in Fig. LABEL:fig:pl-layers, which shows that stacking more PL blocks consistently improves accuracy. The accuracy saturates around 5 blocks and there is still non-negligeble improvement from 3 $\rightarrow$ 4 blocks. Note that we use $3$ blocks in our COCO and Pascal VOC experiments and $5$ blocks in LVIS ones.
Background classes set $\mathcal{B}$. We study the impacts of using different background classes sets $\mathcal{B}$ in Tab. LABEL:tab:abla-B. The first and second rows represent using the stuff (non-object) classes, e.g., floor, sky, of COCOstuff [3] and ADE20k [57] to construct $\mathcal{B}$. Note that COCOstuff has 53 stuff classes and ADE20k has 35 stuff classes, which explains the minor accuracy drop from ADE20k. The third row denotes using all classes of ADE20k as $\mathcal{B}$. This indicates that adding thing (object) classes to $\mathcal{B}$ may not be beneficial. Tab. LABEL:tab:abla-B shows that changing $\mathcal{B}$ only results in small accuracy variations. This suggests that our proposed feature subspace projection does not rely on specific prototypes.
Supervised object detection through region propagation. We further study the application of our proposed techniques in supervised object detection. In Fig. LABEL:fig:reg-loss, we compare the mAPs and regression loss curves of a standard RCNN head and using region propagation for RCNN prediction in COCO detection. We use a ResNet-50 backbone and the same number of parameters for each head. Despite having similar mAPs (37.15 vs. 37.23), Fig. LABEL:fig:reg-loss shows that our proposed region propagation produces a consistently lower regression loss than the standard RCNN. This suggests the potential of our proposed techniques in object detection tasks.
We provide further analysis on the effects of the number of shots, scenarios when expanded proposals cover multiple same-class objects, and more qualitative visualizations in the supplementary document.
Table 9: Ablation studies on the region propagation-based localization.